# August 27, 2025

### v5

@@ Abstract
In this short-length paper, we will present some math that play a central role in quantum hydrodynamics and were presented by Mita in the past. In this formulation of QM, a quantity is involved which is the diffusion flux. As we will see, this quantity plays a crucial role on the evolution of wave packets. More specifically, we will initially briefly explain some results derived in Mita’s papers, by analyzing Gaussian and soliton wave packets, while our main focus is on the role of diffusion in the evolution of super-luminal wave packets. Also, a very brief discussion will be made about the presence of the osmotic velocity field in quantum dynamics.
@@ I Introduction
$\overrightarrow{D}(r,t)=\frac{\hbar}{M}Re\left(\Psi^{*}(r,t)\overrightarrow{\nabla}\Psi(r,t)\right)$
@@ II THE OSMOTIC VELOCITY
. The role of that velocity according to the authors was to balance diffusion in stationary state systems, where there was no probability evolution over time. This velocity, according to the authors, is responsible for ”guiding” the particle in higher probability areas. It was also noted that the osmotic velocity becomes infinite at the nodes of the wave function. Based on that, we’ll give our own interpretation of this quantity. To be more specific, we know that for every position square-integrable wave function $\Psi(x)$, we have a corresponding momentum wave function $\Psi(\rho)$ which is also square-integrable. This means that $|\Psi(\rho\rightarrow\infty)|^{2}=0$. In other words, it’s very difficult to find a particle that has extremely large velocity. And indeed, this is what we observe with the osmotic velocity. The particle has very large osmotic velocity close to the nodes of the wave function, where it is obviously very difficult to be observed. And that, of course, sounds very natural since the dwell time of a particle at a specific spatial region is inversely proportional to its velocity in that region. In a similar way, we can explain why the osmotic velocity approaches zero in the highest probability areas. And, of course, taking into advantage of the osmotic velocity expression, Eq. (3) gives us in stationary states that $E_{K,x}^{stationary}=\frac{1}{2}M\int_{\mathbb{R}}\rho(x)u^{2}(x)dx=\frac{1}{2}M<u^{2}(x)>$, as expected.
@@
III DIFFUSION FLUX IN GAUSSIAN AND SOLITON WAVE PACKETS
@@
The first term is related to the ”centre of mass” motion of the wave packet and the second term, which contains the diffusion flux is related to the expansion of the probability distribution. After very long time, the system approaches an equilibrium state in which we observe that the wave packet diffusivity approaches zero. Indeed, let’s choose for simplicity that $u_{0}=0$ and assume that $t>>0$ and $|x|<<\infty$. We observe that since
@@
where $|s|<<1$. Now, when it comes to the soliton wave packet, Mita mita2021schrodinger defined a probability distribution:
@@
$V(x,t)=-\frac{2\hbar^{2}}{m\sigma_{0}}\rho_{sol}(x,t)$
. At first sight, from a hydrodynamical perspective, this equation doesn’t give us any physical intuition. However, this equation can be written in the following form:
$\overrightarrow{F}(x,t)=\frac{2\hbar^{2}}{m\sigma_{0}}\overrightarrow{\nabla}\rho_{sol}(x,t)\propto-\overrightarrow{D}_{sol}(x,t)$
And that’s how the form of the wave packet is preserved. The greater the tendency of the wave packet to diffuse (greater $\overrightarrow{D}$) in a specific direction, the larger the ”effective” force of opposite direction that has to be applied. And, of course, the coefficient $\mu=\frac{2\hbar^{2}}{m\sigma_{0}}$ is position independent, because otherwise it would cause deformation of the wave packet.
@@
IV DIFFUSION AND SUPER-LUMINAL WAVE PACKETS
In this Chapter, we will examine a state that at $t=0$ is extremely localized. More specifically, we wiil consider the initial state:
@@ V DISCUSSIONS AND CONCLUSIONS
In this paper, we managed to give some clues to support the idea of contribution of the osmotic velocity field on quantum dynamics, even on stationary states where Bohm’s/current velocity plays no role. We also presented the mathematical and physical/hydrodynamical reason behind the infinite diffusion of strictly localized wave packets. So, one the one hand, if someone wants to examine the evolution of a wave-packet, they’d better look at the Bohmian current velocity, but on the other hand, as we saw, diffusion flux also provides us with significant information relevant to the origin of quantum diffusion, which arises from the diffusion tendencies of the initial state (that are expressed by the diffusion flux).

### v6

# September 9, 2025

@@ Abstract
In this short-length paper, we will present some math that play a central role in quantum hydrodynamics and were presented by Mita in the past. In this formulation of QM, a quantity is involved which is the diffusion flux. As we will see, this quantity plays a crucial role on the evolution of wave packets. More specifically, we will initially briefly explain some results derived in Mita’s papers, by analyzing Gaussian and soliton wave packets, while we also focus on the role of diffusion flux in the evolution of the square wave packet. Also, a very brief discussion will be made about the possible presence of the osmotic velocity field in quantum dynamics.
@@ I Introduction
$\overrightarrow{D}(r,t)=-\frac{\hbar}{M}Re\left(\Psi^{*}(r,t)\overrightarrow{\nabla}\Psi(r,t)\right)$
@@ II THE OSMOTIC VELOCITY
. The role of that velocity according to the authors was to balance diffusion in stationary state systems, where there was no probability evolution over time. This velocity, according to the authors, is responsible for ”guiding” the particle in higher probability areas. It was also noted that the osmotic velocity becomes infinite at the nodes of the wave function. Based on that, we’ll give our own interpretation of this quantity. To be more specific, we know that for every position square-integrable wave function $\Psi(x)$, we have a corresponding momentum wave function $\Psi(\rho)$ which is also square-integrable. This means that $|\Psi(\rho\rightarrow\infty)|^{2}=0$. In other words, it’s very difficult to find a particle that has extremely large velocity. And indeed, this is what we observe with the osmotic velocity. The particle has very large osmotic velocity close to the nodes of the wave function, where it is obviously very difficult to be observed. And that, of course, sounds very natural since the dwell time of a particle at a specific spatial region is inversely proportional to its velocity in that region (there is no boundary in the velocity since we are talking about non-relativistic QM). In a similar way, we can explain why the osmotic velocity approaches zero in the highest probability areas. And, of course, by taking into advantage the osmotic velocity expression, Eq. (3) gives us in stationary states that $E_{K,x}^{stationary}=\frac{1}{2}M\int_{\mathbb{R}}\rho(x)u^{2}(x)dx=\frac{1}{2}M<u^{2}(x)>$, as expected.
@@
III DIFFUSION FLUX IN GAUSSIAN WAVE PACKETS AND SOLITONS
@@
The first term is related to the ”centre of mass” motion of the wave packet and the second term, which contains the diffusion flux is related to the expansion of the probability distribution. After very long time, the system approaches an equilibrium state in which we observe that the wave packet diffusive tendency approaches zero. Indeed, let’s choose for simplicity that $u_{0}=0$ and assume that $t>>0$ and $|x|<<\infty$. We observe that since
@@
where $|s|<<1$. Also, $|\overrightarrow{D}_{g}(x,t=t_{2}>>0)|<|\overrightarrow{D}_{g}(x,t=t_{1}>>0)|$, where $t_{1}<t_{2}$ implying that the diffusion tendency decreases with time for large enough times. Now, when it comes to the soliton wave packet, Mita mita2021schrodinger defined a probability distribution:
@@
$V(x,t)=-\frac{2\hbar^{2}}{m\sigma_{0}}\rho_{sol}(x,t)$ (4)
At first sight, from a hydrodynamical perspective, this equation doesn’t give us any physical intuition. However, what we will do is to write the equation in the following form:
$\overrightarrow{F}(x,t)=\frac{2\hbar^{2}}{m\sigma_{0}}\overrightarrow{\nabla}\rho_{sol}(x,t)\propto-\overrightarrow{D}_{sol}(x,t)$ (5)
And that’s how the form of the wave packet is preserved. The greater the tendency of the wave packet to diffuse (greater $\overrightarrow{D}$) in a specific direction, the larger the ”effective” force of opposite direction that has to be applied. And, of course, the coefficient $\mu=\frac{2\hbar^{2}}{m\sigma_{0}}$ is position-independent, because otherwise it would cause deformation of the wave packet.
@@
It must be remarked that the Eq. (5) appears to be the fundamental condition for production of soliton solutions rather than Eq. (4) since the transformation $V^{\prime}(x,t)=V(x,t)+\gamma(t)$, under which Eq. (5) remains invariant, produces solutions of the form:
$\Psi^{\prime}_{sol}(x,t)=\Psi_{sol}(x,t)e^{-i\gamma(t)/\hbar}$
that behave like solitons too.
IV DIFFUSION AND SQUARE WAVE PACKETS
In this Chapter, we will examine a state that at $t=0$ is extremely localized. More specifically, we will consider the initial state:
@@ V DISCUSSIONS AND CONCLUSIONS
In this paper, we managed to give some clues to support the idea of contribution of the osmotic velocity field on quantum dynamics, even on stationary states where Bohm’s/current velocity plays no role. We discussed the reasons behind the relaxation to equilibrium of Gaussian wave packets after very long time and explained solitons’s behavior as a result of counterbalance between diffusion flux and effective forces. We also presented the mathematical and physical/hydrodynamical reason behind the infinite diffusion of strictly localized square wave packets. So, one the one hand, if someone wants to examine the evolution of a wave-packet, they’d better look at the Bohmian current velocity, but on the other hand, as we saw, diffusion flux also provides us with significant information relevant to the origin of quantum diffusion, which arises from the diffusion tendencies of the initial state (that are expressed by the diffusion flux).